How do you turn the equation h=-5t^2+20t+105 into vertex form? Thank you
Answer by ewatrrr(24785)

 
Hi,
Finding vertex form of:
h = -5t^2+20t+105 |Completing the Square
h = -5(t^2 - 4t) + 105 |Note: t^2 -4t + 4 = (t-2)^2
h= -5(t-2)^2 + 20 + 105
h= -5(t-2)^2 + 125 V(2,125)